# A variational approach to dimension-free self-normalized concentration

### v1

@@ 4.1 Optimizing $\lambda$ via Stitching
We begin with sub-gamma processes. Fortunately, the arithmetic is sufficiently navigable in this case so as to allow us to give a bound with small and precise constants. The result is Theorem 4.3 below. As we mentioned in Section 2, all sub-$\psi$ functions are sub-gamma, in the sense that for all twice-differentiable $\psi:[0,\lambda_{\max})\to\mathbb{R}$ with $\psi(0)=\lim_{x\to 0+}\psi^{\prime}(x)=0$, there exist constants $a,c>0$ such that $\psi(\lambda)\leq a\psi_{G,c}(\lambda)$ [33, Proposition 1]. Sub-gamma bounds can thus be applied to all sub-$\psi$ processes after appropriate scaling.
As far as stitching is concerned, in the proof of Theorem 4.3 the epochs are defined relative to the intrinsic time $\det V_{t}$. More specifically, the $k$-th epoch is $E_{k}=\{t:\eta^{k}\leq\det V_{t}<\eta^{k+1}\}$ for some parameter $\eta>1$. The weight assigned to the $k$-th epoch is determined by a “stitching function” $\ell(k)$, which obeys $\sum_{k}\ell^{-1}(k)\leq 1$. In the result that follows we make the simplifying choices $\eta=2$ and $\ell(k)=(k+1)^{2}\zeta(2)$ where $\zeta(2)=\pi^{2}/6$ is the Riemann-zeta function. A more general theorem which keeps all hyperparameters unspecified—Theorem B.1—from which this result follows can be found in Appendix B.3.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi_{G,c}$ process for any $c>0$ and suppose that $V_{t}\succeq U_{0}\succeq I_{d}$ for all $t$ and some positive-definite $U_{0}$ with $\rho=\gamma_{\min}(U_{0})$. Fix $\delta\in(0,1/\sqrt{2}]$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{cD_{\tau}+1.60\sqrt{D_{\tau}}+\max\left\{\frac{c+\sqrt{c^{2}+2\rho}}{2\rho},\sqrt{\frac{D_{\tau}}{2}}\right\}}{H_{c,\tau}},$ (4.5)
@@
$D_{\tau}=\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right)+1.5+2\log(\log_{2}(\det V_{\tau})+1)+\log(1/\delta),$ (4.6)
@@
$H_{c,\tau}=0\vee\left(\sqrt{1-\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}\frac{2}{(c+\sqrt{c^{2}+2c})}}\right).$ (4.7)
If the maximum in (4.5) is achieved by $\sqrt{D_{t}/2}$, then we can replace the numerator with $cD_{\tau}+2.3\sqrt{D_{\tau}}$. The denominator should be viewed as a penalty we pay for the appearance of $g_{\psi,t}$ in Theorem 4.1. If $\gamma_{\min}(V_{t})\to\infty$ then $H_{c,t}\to 1$, just as $g_{\psi,t}$. We suspect that the appearance of $H_{c,\tau}$ in (4.5) is sub-optimal, but it’s unclear how to remove this term with our current proof techniques.
Theorem 4.3 assumes that $V_{t}\succeq I_{d}$ to ensure that the union of the epochs $E_{k}=\{t:\eta^{k}\leq\det V_{t}<\eta^{k+1}\}$, $k\geq 0$ used in the stitching argument constitutes a partitioning of the sample space. If this assumption is not met, however, we can rescale the result as follows. If $V_{t}\succeq\beta I_{d}$ for some $\beta<1$ and $(S_{t},V_{t})$ is a sub-$\psi$ process then Whitehouse et al. [60, Proposition 2.4] shows that $(S_{t}/\sqrt{\beta},V_{t}/\beta)$ is a sub-$\psi_{\beta}$ process for $\psi_{\beta}(\lambda)=\beta\psi(\lambda/\sqrt{\beta})$. We can thus apply Theorem 4.3 to this rescaled process.
@@
The appearance of both the logarithmic term $\log\det V_{\tau}$ and the iterated logarithmic term $\log\log\det V_{\tau}$ in Theorem 4.3 may seem strange. They have different origins: the latter is a result of stitching while the former comes from the variational proof technique and, in particular, is the KL-divergence between the prior and posterior in Proposition 2.2. In short, the term $\log\det V_{\tau}$ captures the complexity of the process. It replaces $d\log\kappa(V_{\tau})$ in the bounds of Whitehouse et al. [60] and is comparable to the dependence on variance-like terms in non self-normalized bounds (see, e.g., [13, Theorem 2.3]).
@@
Related to the previous remark, it’s worth noting that the appearance of $\log\det V_{\tau}$ has unfortunate consequences when the results are applied in $d=1$. In this case, $\log\kappa(V_{\tau})=0$ and the bounds of Whitehouse et al. [60] scale as $\|S_{\tau}\|_{2}\lesssim\sqrt{V_{\tau}}(\psi^{*})^{-1}(\log\log(V_{\tau})+\log(1/\delta))$. In our case, however, we retain a $\log(V_{\tau})$ term, which is unusual for time-uniform bounds in the scalar setting (cf. [34]). In particular, one might expect the term $\log\log(V_{\tau})$ instead of $\log(V_{\tau})$. In the multivariate setting, however, an iterated logarithm dependence on $\det V_{\tau}$ is impossible, as shown by the following example.
@@
We borrow the conclusion of an elegant example from de la Peña et al. [15], also studied by Whitehouse et al. [60] (see their Example 4.6). In particular, de la Peña et al. [15] construct a sub-Gaussian process $(S_{t},V_{t})$ in $\mathbb{R}^{2}$ such that $\|S_{t}\|_{V_{t}^{-1}}\sim\log(t)$, $\gamma_{\max}(V_{t})\sim t(1+a)$, and $\gamma_{\min}(V_{t})\sim\log(t)/(1+)$ for a constant $a$. Hence $\det V_{t}\sim t\log t$ implying that $\|S_{t}\|_{V_{t}^{-1}}\sim\log(\det V_{t})$.
(a) Small $(\det V_{\tau})$
(b) Moderate $\det(V_{\tau})$
(c) Large $\det(V_{\tau})$
Figure 2: A comparison of Theorem 4.3 with the bound of Whitehouse et al. [60], which we recall is based on the condition number of $V_{\tau}$. We control the growth of the determinant with rank-$k$ updates: as $k$ grows, so does $\det V_{\tau}$. The condition number has the same growth in each case. Theorem 4.3 outperforms the condition number-based bound for smaller values of $\det V_{\tau}$ and loses ground as the determinant grows relative to $d\log\kappa(V_{\tau})$. We use $d=20,c=1$ and $U_{0}=I_{d}$. Full simulation details can be found in Appendix C.
Figure 2 compares Theorem 4.3 to the sub-gamma bound presented by Whitehouse et al. [59]. We fix the growth rate of the condition number, $\kappa(V_{t})$, and vary that of $\det(V_{t})$. As expected, our bound outperforms the condition number-based bound of Whitehouse et al. [60] when the determinant grows slowly but is uniformly weaker when the determinant grows quickly.
Intriguingly, the bound of Whitehouse et al. [60] seems to always dominate our bound at small sample sizes (roughly <200<200). The is due to the factor of $H_{c,t}$ in our bounds, which is small at low sample sizes (or, more specifically, until $\gamma_{\min}(V_{t})$ grows sufficiently large). This highlights a drawback of our bounds: if $\gamma_{\min}(V_{t})$ never grows, or grows extremely slowly, then $H_{c,t}$ can remain extremely small, blowing up the bound and making it vacuous.
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}\lesssim\frac{(\psi_{G,c}^{*})^{-1}(D_{\tau})}{H_{c,\tau}}.$ (4.8)
The appearance of $(\psi_{G,c}^{*})^{-1}$ is comforting: reliance on the inverse convex conjugate of $\psi$ is consistent with past work on sub-$\psi$ concentration, such as Whitehouse et al. [60] (see (1.1) and (1.2) in Section 1.1) and Manole and Ramdas [41]. The reliance here also suggests that we may be able to obtain a bound for general $\psi$ that relies on $(\psi^{*})^{-1}$. This is the goal of Theorem 4.7 below.
Before we get there, however, let us compare Theorem 4.3 with the line-crossing inequality in Theorem 4.1. Figure 3 examines the sub-$\psi_{G,c}$ boundary both for fixed values of $\lambda$ and when we stitch over $\lambda$. For reasonable sample sizes, some values of $\lambda$ dominate the stitched boundary of Theorem 4.3 (dotted black line). This is not out of step with previous work: stitching is often deployed as a theoretical tool to obtain optimal rates, not necessarily as a practical bound. We find that smaller values of $\lambda$ tend to do better, but not monotonically. In practice we recommend sample splitting (if possible) to choose $\lambda$.
Let us now move on to Theorem 4.7, which explicitly relates the growth of $\|S_{t}\|_{V_{t}^{-1}}$ to the inverse of $\psi^{*}$. As stated, the result is more of mathematical interest than of practical value. We consider general CGF-like functions $\psi$ which have a nonnegative third derivative and $\lambda_{\max}<\infty$, and show how to relate the growth rate to the convex conjugate of $\psi$. Since we stitch over epochs of the form $\{t:\eta^{k}\leq\|S_{\tau}\|_{V_{\tau}^{-1}}<\eta^{k+1}\}$, the iterated logarithm term of the bound is a function of $\|S_{\tau}\|_{V_{\tau}^{-1}}$ itself instead of $\det V_{\tau}$ as in Theorem 4.3. Such self-reference makes the bound difficult to apply (though this is remedied somewhat by Corollary 4.8 below). Still, as iterated logarithm terms grow extremely slowly, [^11] For instance, $\log_{2}(\log_{2}(10^{200}))<10$. $10^{200}$ is roughly the number of books in Jorge Luis Borges’ Library of Babel [8], while the approximate number of the atoms in the observable universe is a meager $10^{80}$. we feel the result gives a sense of the growth rate of $\|S_{t}\|_{V_{t}^{-1}}$. The proof is provided in Appendix B.4. We refer again to Remark 4.2 which points out that the assumption that $\psi(\lambda)/\lambda^{2}$ is nondecreasing is a minor one.
Figure 3: Comparison of our bound instantiated with a fixed $\lambda$ (Theorem 4.1) versus stitching over lambda (Theorem 4.3). Here $\psi=\psi_{G,c}$ with $c=1/4$. Darker green lines correspond to higher values of $\lambda$.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi$ process where $\psi$ is CGF-like, $\psi(\lambda)/\lambda^{2}$ is nondecreasing, and $\psi^{\prime\prime\prime}=\frac{\textnormal{d}^{3}\psi}{\textnormal{d}\lambda^{3}}\geq 0$. In addition, assume that $\lambda_{\max}<\infty$ and suppose that for all $t\geq 1$, $V_{t}\succeq U_{0}$ where $U_{0}$ is positive-definite and satisfies $\gamma_{\min}^{-1}(U_{0})<\psi(\lambda_{\max})$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\lesssim(\psi^{*})^{-1}\left(\frac{\log(\det V_{\tau})+\log(\log(\|S_{\tau}\|_{V_{\tau}^{-1}}))+\log(1/\delta)}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}\right).$ (4.9)
If we are prepared to sacrifice some tightness, we can move the iterated logarithm term outside of the convex conjugate in (4.9) and bound it by a constant times $\|S_{\tau}\|_{V_{\tau}^{-1}}$. This results in the following bound. The multiplicative factor in front of the convex conjugate is the cost of removing the self-reference from Theorem 4.7. The proof is in Appendix B.5.
@@
Consider the same setup as Theorem 4.7. Let $\varphi(y)=(\psi^{*})^{-1}(y)$ and $\rho=\gamma_{\min}(U_{0})$. Fix $\delta\in(0,1)$ and suppose that $\varphi^{\prime}(\log(1/\delta))<\sqrt{1-1/\rho}$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\lesssim\left(\frac{\sqrt{1-1/\rho}}{\sqrt{1-1/\rho}-\varphi^{\prime}(\log(1/\delta))}\right)(\psi^{*})^{-1}\left(\frac{\log(\det V_{\tau})+\log(1/\delta)}{\sqrt{1-1/\gamma_{\min}(V_{\tau})}}\right).$ (4.10)
5 Bernstein and Bennett Inequalities
In the scalar settings, two standard inequalities for light-tailed random variables are Bennett’s inequality and Bernstein’s inequality (see Boucheron et al. [9, Theorem 2.9 and Theorem 2.10]). The former holds for bounded random variables and the latter holds under a moment condition which has come to be known as “Bernstein’s condition.” Both results rely on upper bounding the log-MGF with a particular $\psi$-function; $\psi_{P,c}$ in the case of Bennett’s inequality and $\psi_{G,c}$ in the case of Bernstein’s. Here we provide dimension-free, self-normalized versions of these inequalities.

### v2

@@ 4.2 Optimizing $\lambda$ via Stitching
We begin with sub-gamma processes. Fortunately, the arithmetic is sufficiently navigable in this case so as to allow us to give a bound with small and precise constants. The result is Theorem 4.4 below. As we mentioned in Section 2, all sub-$\psi$ functions are sub-gamma, in the sense that for all twice-differentiable $\psi:[0,\lambda_{\max})\to\mathbb{R}$ with $\psi(0)=\lim_{x\to 0+}\psi^{\prime}(x)=0$, there exist constants $a,c>0$ such that $\psi(\lambda)\leq a\psi_{G,c}(\lambda)$ [34, Proposition 1]. Sub-gamma bounds can thus be applied to all sub-$\psi$ processes after appropriate scaling.
As far as stitching is concerned, in the proof of Theorem 4.4 the epochs are defined relative to the intrinsic time $\det V_{t}$. More specifically, the $k$-th epoch is $E_{k}=\{t:\eta^{k}\leq\det V_{t}<\eta^{k+1}\}$ for some parameter $\eta>1$. The weight assigned to the $k$-th epoch is determined by a “stitching function” $\ell(k)$, which obeys $\sum_{k}\ell^{-1}(k)\leq 1$. In the result that follows we make the simplifying choices $\eta=2$ and $\ell(k)=(k+1)^{2}\zeta(2)$ where $\zeta(2)=\pi^{2}/6$ is the Riemann-zeta function. A more general theorem which keeps all hyperparameters unspecified can be found in Appendix B.3.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi_{G,c}$ process for any $c>0$. Further suppose that $V_{1}+U_{0}\succeq I_{d}$ for all $t\geq 1$ and some positive-definite $U_{0}$ with $\upsilon=\gamma_{\min}(U_{0})$. Fix $\delta\in(0,1/e]$. Let $\xi(x)=2/(c+\sqrt{c^{2}+2x})$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{cD_{\tau}+1.51\sqrt{D_{\tau}}+\xi(c)\max\left\{\frac{1}{\xi(\upsilon)\upsilon},\sqrt{\frac{D_{\tau}}{2}}\right\}}{H_{c,\tau}},$ (4.5)
@@
$D_{\tau}=\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1.5+2\log(\log_{2}(\det(V_{\tau}+U_{0}))+1)+\log(1/\delta),$ (4.6)
@@
$H_{c,\tau}=0\vee\left(\sqrt{\frac{\sqrt{c^{2}+2c}-c}{3+c}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t}+U_{0})}\xi(c)}\right).$ (4.7)
The denominator in Theorem B.1 should be viewed as a penalty we pay for the appearance of $g_{\psi,t}$ in Theorem 4.1. If $\gamma_{\min}(V_{t}+U_{0})\to\infty$ then $H_{c,t}$ converges to a constant depending on $c$, just as $g_{\psi,t}$. Note that $H_{c,t}$ is larger than 0 as soon as
$\gamma_{\min}(V_{t}+U_{0})>\frac{(3+c)\gamma_{\max}(U_{0})}{c}.$
Thus, as $c\to 0$, it takes either more observations or a faster growth rate of $\gamma_{\min}(V_{t}+U_{0}$) for the bound to be informative. We suspect that the appearance of $H_{c,\tau}$ in (4.5) is sub-optimal, but it’s unclear how to remove this term with our current proof techniques.
Theorem 4.4 assumes that $V_{1}+U_{0}\succeq I_{d}$ to ensure that the union of the epochs $E_{k}=\{t:\eta^{k}\leq\det(V_{t}+U_{0})<\eta^{k+1}\}$, $k\geq 0$ used in the stitching argument constitutes a partitioning of the sample space. If this assumption is not met, however, we can rescale the result as follows. If $V_{t}\succeq\beta I_{d}$ for some $\beta<1$ and $(S_{t},V_{t})$ is a sub-$\psi$ process then Whitehouse et al. [63, Proposition 2.4] shows that $(S_{t}/\sqrt{\beta},V_{t}/\beta)$ is a sub-$\psi_{\beta}$ process for $\psi_{\beta}(\lambda)=\beta\psi(\lambda/\sqrt{\beta})$. We can thus apply Theorem 4.4 to this rescaled process.
@@
The appearance of both the logarithmic term $\log\det V_{\tau}$ and the iterated logarithmic term $\log\log\det V_{\tau}$ in Theorem 4.4 may seem strange. They have different origins: the latter is a result of stitching while the former comes from the variational proof technique and, in particular, is the KL-divergence between the prior and posterior in Proposition 2.2. In short, the term $\log\det V_{\tau}$ captures the complexity of the process. It replaces $d\log\kappa(V_{\tau})$ in the bounds of Whitehouse et al. [63] and is comparable to the dependence on variance-like terms in non self-normalized bounds (see, e.g., Chugg et al. [15, Theorem 2.3]).
@@
Related to the previous remark, it’s worth noting that the appearance of $\log\det V_{\tau}$ has unfortunate consequences when the results are applied in $d=1$. In this case, $\log\kappa(V_{\tau})=0$ and the bounds of Whitehouse et al. [63] scale as $\|S_{\tau}\|_{2}\lesssim\sqrt{V_{\tau}}(\psi^{*})^{-1}(\log\log(V_{\tau})+\log(1/\delta))$. In our case, however, we retain a $\log(V_{\tau})$ term, which is unusual for time-uniform bounds in the scalar setting (cf. [35]). In particular, one might expect the term $\log\log(V_{\tau})$ instead of $\log(V_{\tau})$. In the multivariate setting, however, an iterated logarithm dependence on $\det V_{\tau}$ only is impossible, as shown by the following example.
@@
We borrow the conclusion of an elegant example from de la Peña et al. [19], also studied by Whitehouse et al. [63] (see their Example 4.6). In particular, de la Peña et al. [19] construct a sub-Gaussian process $(S_{t},V_{t})$ in $\mathbb{R}^{2}$ such that $\|S_{t}\|_{V_{t}^{-1}}\sim\log(t)$, $\gamma_{\max}(V_{t})\sim t(1+a)$, and $\gamma_{\min}(V_{t})\sim\log(t)/(1+a)$ for a constant $a$. Hence $\det V_{t}\sim t\log t$ implying that $\|S_{t}\|_{V_{t}^{-1}}\sim\log(\det V_{t})$.
We also briefly note that Example 4.7 shows that, even in the fixed-time setting, one cannot hope for bounds that do not have logarithmic dependence on $V_{t}$. This is in contrast to the non-self-normalized fixed-time setting, where such dependence can be avoided.
4.3 A note on suspected suboptimality
Consider Theorem 4.4 with $V_{t}$ such that $\gamma_{\min}(V_{t})\to\infty$ and $c=1$. With probability $1-\delta$, we obtain that
$\|S_{t}\|_{(V_{t}+U_{0})^{-1}}\lesssim\log\det(V_{t})+\log(1/\delta)+\sqrt{\log\det(V_{t})+\log(1/\delta)}.$
To put this in more familiar terms, suppose that $V_{t}=tI_{d}$. Then we obtain the bound
$\left\|\frac{S_{t}}{t}\right\|\lesssim\frac{d\log(t)+\log(1/\delta)}{\sqrt{t}}+\sqrt{\frac{d\log(t)+\log(1/\delta)}{t}}.$
This resembles the terms in a Bernstein bound, except for a missing factor of $1/\sqrt{t}$ on the first term. Indeed, we expect the linear term in a Bernstein bound to shrink at rate $1/t$, not $1/\sqrt{t}$, thus making it asymptotically smaller than the square root term. This makes us suspect that Theorem 4.4 is suboptimal in terms of dependence on $D_{t}$. We imagine that the correct dependence would resemble $cD_{t}/\gamma_{\min}(V_{t})+\sqrt{D_{t}}$, akin to the bound of Whitehouse et al. [63] (though we emphasize that theirs is not a log-determinant bound, and thus not dimension-free).
It is instructive to compare our bound with several concurrent works which study log-determinant Bernstein bounds [48, 4, 43]. Like us, Akhavan et al. [4] obtain a bound which scales linearly in $\log\det(V_{t})$. Meanwhile, both Metelli et al. [48] and Martinez-Taboada et al. [43] avoid this linear dependence, but at the cost of a multiplicative dependence between $\log\det(V_{t})$ and $\log(1/\delta)$. That is, their bounds scale as
$\left\|S_{t}\right\|_{(V_{t}+U_{0})^{-1}}\lesssim\log(1/\delta)+\sqrt{\log(\det(V_{t}))\log(1/\delta)}.$
Our bound will thus be tighter for $\log(\det(V_{t}))\lesssim\log(1/\delta)$, and theirs when $\log(1/\delta)\lesssim\log(\det(V_{t}))$. There are other differences as well. For instance, the result of Metelli et al. [48] only applies to bounded observations; those of Martinez-Taboada and Ramdas [45] are not time-uniform or either require a priori upper bounds on $\log\det(V_{t})$. Whether one can obtain a time-uniform log-determinant bound for sub-gamma processes with (i) additive dependence between $\log\det(V_{t})$ and $\log(1/\delta)$ and (ii) a linear term vanishing at a faster rate, is thus an open question.
4.4 A general bound via the convex conjugate
While Theorem 4.4 removes the need to choose $\lambda$, we find that in practice fixing $\lambda$ and applying Theorem 4.1 results in the tightest bounds. Indeed, Theorem 4.1 typically outperforms Theorem 4.4 by a factor of three to five. This is not out of step with previous work: stitching is often deployed as a theoretical tool to obtain optimal rates, not necessarily as a practical bound. If one is interested in using the line-crossing inequality, we find that a good rule of thumb is to choose $\lambda\approx 0.85\lambda_{\max}$. If possible, we recommend sample splitting to choose $\lambda$.
@@ 4.4 A general bound via the convex conjugate
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\lesssim\frac{(\psi_{G,c}^{*})^{-1}(D_{\tau})}{H_{c,\tau}}.$ (4.8)
The appearance of $(\psi_{G,c}^{*})^{-1}$ is comforting: reliance on the inverse convex conjugate of $\psi$ is consistent with past work on sub-$\psi$ concentration, such as Whitehouse et al. [63] (see (1.1) and (1.2) in Section 1.1) and Manole and Ramdas [42]. The reliance here also suggests that we may be able to obtain a bound for general $\psi$ that relies on $(\psi^{*})^{-1}$. This is the goal of Theorem 4.9 below, which explicitly relates the growth of $\|S_{t}\|_{(V_{t}+U_{0})^{-1}}$ to the inverse of $\psi^{*}$.
As stated, Theorem 4.9 is more of mathematical interest than of practical value. We consider general CGF-like functions $\psi$ which have a nonnegative third derivative and show how to relate the growth rate to the convex conjugate of $\psi$. Since we stitch over epochs of the form $\{t:\eta^{k}\leq\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}<\eta^{k+1}\}$, the width of the bound is itself function of $\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}$ itself instead of $\det V_{\tau}$ as in Theorem 4.4. In particular, the width depends on an iterated-logarithm term involving $\|S_{\tau}\|_{(V_{\tau}+U_{0})^{1}}$ and a second term $f_{\psi}(2\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}})$, for a function $f$ defined below. Such self-reference makes the bound difficult to apply. Still, as iterated logarithm terms grow extremely slowly, [^5] For instance, $\log_{2}(\log_{2}(10^{200}))<10$. $10^{200}$ is roughly the number of books in Jorge Luis Borges’ Library of Babel [9], while the approximate number of the atoms in the observable universe is a meager $10^{80}$. and $f$ is often sublinear, we feel the result gives a sense of the growth rate of $\|S_{t}\|_{(V_{t}+U_{0})^{-1}}$.
To state the result, for a given CGF-like $\psi$, let
$\lambda^{*}(a):=\operatorname*{\textnormal{argmax}}_{\lambda\in(0,\lambda_{\max})}\lambda a-\psi(\lambda),$ (4.9)
which we note is well-defined; see Lemma B.3 in the appendix. Then define the function $f_{\psi}:\mathbb{R}\to\mathbb{R}$ as
$f_{\psi}(x):=\psi(\lambda^{*}(x))\left(\sqrt{\psi^{-1}(\lambda^{*}(x))}-1\right).$ (4.10)
The function $f_{\psi}$ is somewhat opaque, so let us give a few examples of how it grows as a function of $a$. The proof is provided in Appendix B.4.
Lemma 4.8.
Let $(a_{n})$ be an increasing sequence. If $\psi=\psi_{G,c}$ for some $c>0$, then
$f_{\psi}(a_{n})=O\left(\frac{\sqrt{a_{n}}}{c^{3/2}}\right)\left(\sqrt{\frac{2}{c+\sqrt{c^{2}+2c}}}-1\right).$ (4.11)
If $\psi=\psi_{E,c}$ for some $c>0$, then
$f_{\psi}(a_{n})=O(\log(ca_{n}))\left(\frac{1}{\sqrt{c}}-1\right).$ (4.12)
Here the asymptotics are in terms of the sequence $(a_{n})$. We’ve dropped other constants but kept the dependence on $c$. Note that in both cases, if $c$ is large enough, then $f(a_{n})$ is negative.
We now state Theorem 4.9. The proof is provided in Appendix B.5. We refer again to Remark 4.2 which points out that the assumption that $\psi(\lambda)/\lambda^{2}$ is nondecreasing is a minor one.
@@
Let $(S_{t},V_{t})$ be a sub-$\psi$ process where $\psi$ is CGF-like, $\psi(\lambda)/\lambda^{2}$ is nondecreasing, and $\psi^{\prime\prime\prime}=\frac{\textnormal{d}^{3}\psi}{\textnormal{d}\lambda^{3}}\geq 0$. Let $U_{0}$ be positive-definite and satisfy $\gamma_{\min}^{-1}(U_{0})<\psi(\lambda_{\max})$. Set $\widehat{V}_{t}=V_{t}+U_{0}$ for all $t$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\lesssim(\psi^{*})^{-1}\left(\frac{\log(\det\widehat{V}_{\tau})+\log(\log(\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}})/\delta)}{0\vee\sqrt{1-\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{\tau})}}+f_{\psi}\left(2\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\right)\right).$ (4.13)
Given that $f_{\psi}(u)$ scales as $\sqrt{u}$ for $\psi=\psi_{G,c}$ and as $\log(u)$ for $\psi=\psi_{E,c}$, a natural question is whether one should convert a given sub-gamma process into a sub-exponential process before applying Theorem 4.9. Indeed, any sub-$\psi_{G,c}$ is also a sub-$\psi_{E,3c/2}$ process (see Howard et al. [34, Table 5]). Such a transformation will not make much of a difference, however. Both $(\psi_{G,c}^{*})^{-1}$ and $(\psi_{E,c}^{*})^{-1}$ are dominated by linear terms (indeed $(\psi_{G,c}^{*})^{-1}(u)=cu+\sqrt{2u}$ and $(\psi_{E,c}^{*})^{-1}(u)=cu+O(c^{-1}\log(1+c^{2}u))$). Since $f_{\psi}$ is sublinear in both cases, if we write out the bounds and bring all the terms depending on $\|S_{t}\|$ to one side, the difference between the sub-gamma and sub-exponential cases is a bound on $\|S_{t}\|-O(\sqrt{\|S_{t}\|})$ vs $\|S_{t}\|-O(\log(\|S_{t}\|)$, which have the same growth rate.
If we are prepared to sacrifice some tightness and to assume that $f$ is sublinear (as is the case in Lemma 4.8) we can move the terms depending on $\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}$ outside of the convex conjugate in (4.13) and bound it by a constant times $\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}$. This results in Corollary 4.10. The proof is in Appendix B.6.
@@
Consider the same setup as Theorem 4.9. Let $\varphi(y)=(\psi^{*})^{-1}(y)$ and $\upsilon=\gamma_{\min}(U_{0})>1$. For any $0<b<1$, define the hitting time $\tau_{0}=\inf\{t:\upsilon/\gamma_{\min}(\widehat{V}_{t})\leq b\}$. Fix $\delta\in(0,1)$ and suppose both that $f_{\psi}(x)\leq\kappa x$ for some $\kappa>0$ and all $x\geq 0$, and $\varphi^{\prime}(\log(1/\delta))(1+2\kappa)<\sqrt{1-b}$. Then, with probability $1-\delta$, for all stopping times $\tau\geq\tau_{0}$,
$\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}\lesssim\left(\frac{\sqrt{1-b}}{\sqrt{1-b}-\varphi^{\prime}(\log(1/\delta))(1+2\kappa)}\right)(\psi^{*})^{-1}\left(\frac{\log(\det\widehat{V}_{\tau})+\log(1/\delta)}{\sqrt{1-\gamma_{\max}(U_{0})/\gamma_{\min}(\widehat{V}_{\tau})}}\right).$ (4.14)
The hidden constants in (4.14) are those inherited from (4.13). We opt to keep the constant $\nicefrac{{\sqrt{1-b}}}{{\sqrt{1-b}-\varphi^{\prime}(\log(1/\delta))(1+2\kappa)}}$ in the bound so that the cost of removing the reference to $\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}$ on the right hand of (4.13) is made clear.
5 Self-normalized Bernstein and Bennett Inequalities
In the scalar settings, two standard inequalities for light-tailed random variables are Bennett’s inequality and Bernstein’s inequality; see Boucheron et al. [10, Theorem 2.9 and Theorem 2.10]. The former holds for bounded random variables and the latter holds under a moment condition which has come to be known as “Bernstein’s condition.” Both results rely on upper bounding the log-MGF with a particular $\psi$-function; $\psi_{P,c}$ in the case of Bennett’s inequality and $\psi_{G,c}$ in the case of Bernstein’s. Here we provide dimension-free, self-normalized versions of these inequalities.